OpenStudy (anonymous):

cos^2 x csc^2 x + cos^2 x sec^2 x =csc^2 x determine whether the equation is an identity, if it is, prove it..

OpenStudy (twopointinfinity):

oh my.. @mathmale Do you know trig?

OpenStudy (mathmale):

Yes. However, I've been on the Internet most of the morning and need to get off. Hope you can find someone else to guide you along the way towards proving this identity. Good luck!

OpenStudy (anonymous):

Left side of the equation: cos^2 x(1/sin^2 x) + cos^2 x*(1/cos^2 x) = cos^2 x/sin^2 x + 1 = cos^2 x /sin^2x + sin^2 x/sin^2 x = (cos ^2 x + sin^2 x)/sin^2 x = 1/sin^2 x = csc^2 x. It is an identity.